OpenStudy (anonymous):

Medal to first right answer!! NEED HELP QUICK! For the function c(x) = x^2 +4x -5/ x+5. a. Identify the restricted domain value(s). b. Simplify the function. c. Graph the function. d. Make a table of values.

OpenStudy (anonymous):

Need help with letter C

OpenStudy (anonymous):

The domain value is \[x \neq5\]

OpenStudy (anonymous):

the domain value \[x \neq -5\]

OpenStudy (anonymous):

letter B is \[c(x)=(x+5)\times (x-1)/ (x+5) C(x)=x-1\]

OpenStudy (anonymous):

Oh sorry didn't see that i missed the negative sign! Thanks!

OpenStudy (anonymous):

I have letter C (the table of values) and will give if it helps

OpenStudy (anonymous):

Can u help me with Letter D Abdulhameed?

OpenStudy (anonymous):

oops i mean C

OpenStudy (anonymous):

hi i'm new in this website i've just signed up :) , ok i will try

OpenStudy (anonymous):

x^2 +4x -5/ x+5=f(x) the easy way to do this is to write the equation in vertex form first. vertex form is the form f(x)=a(x-h)^2+k where h, k is the vertex (the highest or lowest form) of the parabola.

OpenStudy (anonymous):

after that you pick points on the parabola and connect them with a smooth line.

OpenStudy (anonymous):

@yolo497 do u want only the graph or the steps to the graph ??

OpenStudy (anonymous):

Both would be nice, so i could do other problems in the future by myself

OpenStudy (anonymous):

Thanks @Abdulhameed

OpenStudy (anonymous):

I will leave it to @Abdulhameed

OpenStudy (anonymous):

No im listening to both of you

OpenStudy (anonymous):

go ahead im following along

OpenStudy (anonymous):

Trust me, it's better if one person answers. If you don't get his answer I can talk after.

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

@Abdulhameed maybe u could give me the graph first and then teach me how to do it?

OpenStudy (anonymous):

if that's easier?

OpenStudy (anonymous):

is it 4x-5/x+5 or (4x-5)/x+5 ???

OpenStudy (anonymous):

4x-5/x+5

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):